# Can Emulating Semantic Translation Help LLMs with Code Translation? A Study Based on Pseudocode

### v2

@@ 6.1. Automated Code Translation
To facilitate development activities relying on code translation, various automated code translation methods have emerged in the past decades (Chen et al., 2025; Jiao et al., 2023). Nguyen et al. (2013) pioneered in automated code translation via statistical machine translation. To improve translation accuracy by operating directly on ASTs or program graphs, researchers designed tree- and graph-based neural models (Chen et al., 2018; Guo et al., 2021). Unsupervised and self-supervised model training approaches were later proposed to mitigate the scarcity of parallel code corpora (Rozière et al., 2020, 2022; Liu et al., 2023).
@@
Existing automated code translation approaches mainly focus on code-to-code transformation. Differently, we study the effectiveness of emulating semantic translation for code via pseudocode, where LLMs first describe the semantics of original programs with pseudocode and then produce translated programs. We identify the complementary role of pseudocode-based translation for conventional direct translation, as well as its helpfulness in reducing task difficulties and bridging the differences between the features and APIs in varying PL pairs. The findings may guide practitioners to obtain accurate LLM-generated code translations by leveraging pseudocode.
In addition, some works attempted to debug the erroneous code translations, e.g., by locating code snippets leading to inconsistent behaviors between the original and translated programs to guide repair (Wang et al., 2023b) and conducting certain types of repair based on fixed templates (Ramos et al., 2024). The methods mainly debug minor mistakes in the translated code. They are generally orthogonal to our pseudocode-based code translation approach, which aims to reduce the occurrence of errors in the translation results and may also facilitate patch generation. Our methods can be combined with the debugging methods to further enhance the accuracy of LLM-generated code translations.
@@ 6.2. Applications of Pseudocode
In addition to code generation, pseudocode has also proven useful in cross-PL code retrieval by embedding structured pseudocode representations (Kulkarni et al., [n. d.]), in binary function similarity and vulnerability search by leveraging decompiled pseudocode to obtain platform-agnostic semantic representations (Zhang et al., 2022), and in generating descriptive summaries for stripped binaries using pseudocode extracted by decompilers as an expert-guided signal (Ye et al., 2023). All these applications leveraged the advantage of pseudocode in succinctly capturing the intent and logic of programs to facilitate code comprehension and semantic representation (Oda et al., 2015).
This work studies the helpfulness of pseudocode for another common software development task, i.e., code translation, where pseudocode serves as a semantic interpretation to facilitate semantic translation. The extensive investigation of pseudocode’s effectiveness in code translation yields various findings, which may guide practitioners to obtain accurate LLM-generated code translations based on pseudocode. We also identified limitations of pseudocode generated by LLMs, which are expected to shed light on the focus of validation and repair for automatically generated pseudocode, helping harness the potential of pseudocode for downstream tasks like code translation.
@@ 7. Conclusion
In this work, we empirically study the effectiveness of pseudocode-based code translation with LLMs, which explicitly emulates the human practice of semantic translation. By investigating the performance of five popular LLMs with pseudocode-based code translation on 9,690 translation tasks across six popular PLs, we reveal that pseudocode-based code translation can effectively complement the widely adopted direct code translation approach for various pairs of source and target PLs. We also demonstrate the effectiveness of pseudocode as a general and effective intermediary, as well as the further potential of translation based on higher-quality pseudocode and the bottleneck in both code understanding and generation. Our case studies further identify the concrete advantages and limitations of pseudocode-based code translation. Based on these findings, we suggest the hybrid use of pseudocode-based and direct code translation to enhance code translation accuracy, as well as discuss future research directions to further unleash the potential of pseudocode in code translation.

### v3

@@ 6.1. Automated Code Translation
To facilitate development activities relying on code translation, various automated code translation methods have emerged in the past decades (Chen et al., 2025a; Jiao et al., 2023). Nguyen et al. (2013) pioneered in automated code translation via statistical machine translation. To improve translation accuracy by operating directly on ASTs or program graphs, researchers designed tree- and graph-based neural models (Chen et al., 2018; Guo et al., 2021). Unsupervised and self-supervised model training approaches were later proposed to mitigate the scarcity of parallel code corpora (Rozière et al., 2020, 2022; Liu et al., 2023).
@@
Existing automated code translation approaches mainly focus on code-to-code transformation. Differently, we study the effectiveness of emulating semantic translation for code via pseudocode, where LLMs first interpret the semantics of original programs with pseudocode and then produce translated programs. We identify the complementary role of pseudocode-based translation for conventional direct translation, as well as its helpfulness in decoupling task burdens and bridging the differences between the features and APIs in varying PL pairs. The findings may guide practitioners to obtain accurate LLM-generated code translations in a new manner through abstraction.
In addition, some works attempted to debug the erroneous code translations, e.g., by locating code snippets leading to inconsistent behaviors between the original and translated programs to guide repair (Wang et al., 2023b) and conducting certain types of repair based on fixed templates (Ramos et al., 2024). These methods mainly debug minor mistakes in translation results. They are generally orthogonal to our pseudocode-based code translation approach, which aims to reduce the occurrence of errors in the translation results and may also facilitate patch generation. Our methods can be combined with the debugging methods to further enhance the accuracy of LLM-generated code translations.
@@ 6.2. Applications of Pseudocode
In addition to code generation, pseudocode has also proven useful in cross-PL code retrieval by embedding structured pseudocode representations (Kulkarni et al., ), in binary function similarity and vulnerability search by leveraging decompiled pseudocode to obtain platform-agnostic semantic representations (Zhang et al., 2022), and in generating descriptive summaries for stripped binaries using pseudocode extracted by decompilers as an expert-guided signal (Ye et al., 2023). All these applications leveraged the advantage of pseudocode in succinctly capturing the intent and logic of programs to facilitate code comprehension and semantic representation (Oda et al., 2015).
This work studies the helpfulness of pseudocode for another common software development task, i.e., code translation, where pseudocode serves as a semantic interpretation for emulating semantic translation. The extensive investigation of pseudocode’s effectiveness in code translation yields various findings, which may guide practitioners to obtain accurate LLM-generated code translations based on pseudocode. We also identified limitations of pseudocode generated by LLMs, which are expected to shed light on the focus of validation and repair for automatically generated pseudocode, helping harness the potential of pseudocode for downstream tasks like code translation.
@@ 7. Conclusion
In this work, we empirically study the effectiveness of pseudocode-based code translation with LLMs, which explicitly emulates the human practice of semantic translation. By investigating the performance of five popular LLMs with pseudocode-based code translation on 9,690 translation tasks across six popular PLs, we reveal that pseudocode-based translation can effectively complement the widely adopted direct translation approach for various pairs of source and target PLs. We also identify the highly benefited PL pairs and demonstrate the effectiveness of pseudocode as a general and effective intermediary, as well as the further potential of translation based on higher-quality pseudocode and the bottleneck in both code understanding and generation. Our case studies further identify the advantages and limitations of pseudocode-based code translation. Based on these findings, we suggest the hybrid use of pseudocode-based and direct translation approaches to enhance code translation accuracy. We also discuss future research directions to further unleash the potential of pseudocode in code translation.
Acknowledgments
We would like to thank the anonymous reviewers for their insightful comments. This work is partially supported by the National Natural Science Foundation of China (Grant No. 92582201), the Hong Kong SAR Research Grant Council/General Research Fund (Ref No. 16210725), and the Hong Kong SAR Research Grant Council/Theme-based Research Scheme (Ref No. T41-517/25-N).